Two identical guitar strings are stretched with the same tension between supports that are not the same distance apart. The fundamental frequency of the higher-pitched string is 380Hz, and the speed of transverse waves in both wires is 200 m/s. How much longer is the lower-pitched string if the beat frequency is 6Hz?
Answer #1

The velocity of the wave in the string is v=f\lambda

\lambda =\frac{v}{f}

  \lambda =\frac{200}{380}

The wavelength of the wave for the first string is \lambda =\frac{200}{380}= 0.526m

The lenght of the strng is l= \frac{\lambda }{2}=0.263m

the frequency of the second string is f' = 380-6 = 374Hz

The wave length is \lambda ' = \frac{200}{374} = 0.535m

the length of the string is l' =   \frac{\lambda ' }{2}= \frac{0.535}{2}=0.267m   

0.267-0.263 = 0.004m is the change in length in the both the strings.